40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 238 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  5100   Wed Aug 3 01:30:04 2011 JenneUpdateLSCAbsolute length of Xarm and Yarm measured

So far, this is just preliminary, because I haven't done full error analysis to determine the error on my measurements.  That will hopefully be done by tomorrow afternoon (so before we start taking off doors).

I find that the length of the Xarm is:  37.5918 meters.

I find that the length of the Yarm is:   37.5425 meters.

I used the mass-kicking technique, as summarized by Kiwamu, and fully described by Alberto.  More words / description to follow with the full error analysis.

  5099   Tue Aug 2 23:14:21 2011 NicoleUpdateSUSCorrected Screw Alignment in TT Suspension Base

The TT suspension base was not able to be securely mounted to the optical table (i.e. mounted with 4 screws)

because the spacing between the screw holes in the base did not have the correct spacing for mounting on a table with a 1 inch pitch.

newholes.jpg

We have carefully removed the suspension from the problematic base. PLEASE BE VERY CAREFUL AROUND THE TABLE NEXT TO THE MC-2 CHAMBER! THE TT SUSPENSION IS RESTING THERE WITHOUT THE BASE! I will reattach it to the base tomorrow morning when I am less tired and more careful!

We measured the base to be about 4.882" x 3.774". The screw hole spacing is about 3.775" and 2.710" respectively. I have changed the diameter of the screw holes from 0.26" to 0.315" and have been able to successfully mount the suspension to the 1 inch pitch table next to the MC-2 chamber.

 

mountedmirrorflash.jpg

Now that the TT suspension can be mounted, I am going to be aligning a 670nm LED laser and balancing the mirror on the TT tomorrow morning. I will be using a beam blocker but please still be careful.

  5098   Tue Aug 2 23:02:48 2011 NicoleUpdateSUSFixed Accelerometer


 

The EM shaker was broken (the input terminals (banana inputs) had snapped off. To fix this, I have mounted two banana input mounting posts to a metal mount that Steve attached to the shaker.

shakerposts.jpgclipstoamplifier.jpg

However, because bananas do not provide a secure connection (they easily fall out), I have made special wires to connect the banana inputs of the shaker to the mounted banana inputs of the amplifier I am using (along with the sine generating function of the HP 3563A signal analyzer). Upon Koji's suggestion, I have made C-shaped clips to attach to the banana post mounts. These clips are made from insulated ring terminals.

 newclips.jpg

Today I tested the  shaker and it works! WOOT! I currently have the shaker attached to the horizontal sliding platform (without the TT suspension on it).

Using a 750mHz signal from the HP 3563A with an amplitude of 500 mV amplified to 0.75V, I have gotten the shaker to displace the platform (without the TT suspension on it) 1 mm.

  5097   Tue Aug 2 19:05:13 2011 kiwamuUpdateIOOpeparation for low power MC : HWP+PBS installed

[Steve / Kiwamu]

  An attenuator, consisting of two HWPs and a PBS, has been installed on the PSL table for the MC low power state.

Those items allow us to reduce the amount of the incident power going into the MC.

We haven't decreased the power yet because we still have to measure the arm lengths.

After we finish the measurement we will go to the low power state.

We have adjusted the polarization after the last HWP using another PBS. Now it is S-polarizing beam.

After the installation of the attenuator the beam axis has changed although we were immediately able to lock the MC with TEM00 mode.

I touched two steering mirrors on the PSL table to get the transmitted power of MC higher. At the moment the transmitted power in MC_TRANS is at about 30000 cnts.

 

The attached picture is the setup of the attenuator on the PSL table.

Attachment 1: attenuator.png
attenuator.png
  5096   Tue Aug 2 17:40:04 2011 JennyUpdatePSLReducing beam intensity incident on photodiode

I am using a PDA255 photodiode to measure the power outputted by the NPRO beam on the PSL table. (I'm going to then use a network analyzer to measure the amplitude response of the PZT to being driven at a range of frequencies. I'll detect the variation in in response to changing the driving frequency using this PDA255.)

The PDA255 has an active area of 0.8mm^2 and a maximum intensity for which the response is linear of 10mW/cm^2. This means that a beam I focus on the PD must have a power less than 0.08 mW (and even less if the spot size is smaller than the window size).

I used a power meter to measure the beam power and found it was 0.381 mW.

The second polarizing beam splitter in the setup transmits most of the beam power, but reflects 0.04 mW (according to the power meter). I'm going to place the photodiode there in the path of the reflected beam.

  5095   Tue Aug 2 16:55:21 2011 ranaUpdateABSLArm length measurement : cavity kick technique

Quote:

I made some attempts to measure the current length of the arm cavities by using the mass-kicking technique.

 Why not just scan the Green laser to measure the arm lengths instead? The FSR of the arm is ~3.75 MHz and so all you have to do is lock the arm green and then sweep the PZT until the resonance is found at 3.75 MHz.

L.png

  5094   Tue Aug 2 16:43:23 2011 jamieUpdateCDSNDS2 server on mafalda restarted for access to new channels

In order to get access to new DQ channels from the NDS2 server, the NDS2 server needs to be told about the new channels and restarted.  The procedure is as follows:

ssh mafalda
cd /users/jzweizig/nds2-mafalda
./build_channel_history
./install_channel_list
pkill nds2
# wait a few seconds for the process to quit and release the server port
./start_nds2

This procedure needs to be run every time new _DQ channels are added.

We need to set this up as a proper service, so the restart procedure is more elegant.

An additional comment from John Z.:

    The --end-gps parameter in ./build_channel_history seems to be causeing
    some trouble. It should work without this parameter, but there is a
    directory with a gps time of 1297900000 (evidently a test for GPS1G)
    that might screw up the channel list generation. So, it appears that
    the end time requires a time for which data already exists. this
    wouldn't seem to be a big deal, but it means that it has to be modified
    by hand before running. I haven't fixed this yet, but I think that I
    can probably pick out the most recent frame and use that as an end-time
    point. I'll see if I can make that work...

  5093   Tue Aug 2 15:41:06 2011 kiwamuUpdateIOOMeasurement of MC spot positions : done

[Suresh / Kiwamu]

 The measurement of the spot positions on the MC mirrors are DONE.

Surprisingly the spot positions are not so different from the ones measured on May.

    Feb 26 2011      May 08 2011 (New !) Aug 2 2011
MC1 pit [mm]   1.6   1.9  1.93
MC2 pit [mm]   6.4   9.0 9.03
MC3 pit [mm]   1.4   2.0 2.01
MC1 yaw [mm]   -1.5   -1.7 -1.72
MC2 yaw [mm]   1.0   0.2 0.178
MC3 yaw [mm]   -1.3   -1.9 -1.87

 

(some notes)
We used Valera's script senseMCdecenter to estimate the spot positions ( see his entry).

It returns so many EPICS error messages and sometime some measured values were missing. So we had to throw away some of the measurements.

Anyways we gave the resultant ASCII file to Valera's matlab file sensmcass.m to get the actual amount of off-centering in milli-meter.

The attached file is the resultant plot from his matlab code.

Attachment 1: MCdecenter.png
MCdecenter.png
  5092   Tue Aug 2 11:52:44 2011 kiwamuSummaryGeneralSchedule

I have updated the 40m public calender.

Main change :

  + The vent starts from 3rd of August

  + Keiko and Anamaria (LSU) come from 13th of August

  5091   Tue Aug 2 11:02:52 2011 JenneUpdateSUSPRM Watchdog tripped

I found PRM watchdog tripped.  It's all better now.

  5090   Tue Aug 2 10:53:03 2011 JenneOmnistructureSAFETYRegular door out of service. Use Control Room Door only!!

The hazardous waste people are moving chemicals around outside our door, and have roped off our regular front door. 

Please go around, and use the control room door to enter and exit.  It is currently unlocked, although I'll lock up when I leave for LIGOX.

  5089   Tue Aug 2 02:35:23 2011 kiwamuUpdateGeneralpreparation of the vent : status and plan

The vent will take place on Wednesday.

Plan for Tuesday :

  (Morning) Preparation of necessary items for the low power MC (Steve / Jamie)

  (Daytime) Measurement of the MC spot positions (Suresh)

  (Daytime) Arm length measurement (Jenne)

  (Nighttime) Locking of the low power MC (Kiwamu / Volunteers)

 

Plan for Wednesday :

  (Early morning) Final checks on the beam axis, all alignments and green light (Steve / Kiwamu / Volunteers )

  (Morning) Start the vent (Steve)

  (daytime-nighttime) Taking care of the Air/Nitrogen cylinders (Everybody !!)

 

Status of the vent preparation :

 

  (not yet) Low power MC

  (ongoing) Measurement of the arm lengths

  (ongoing) Measurement of the MC spot positions

  (80% done) Estimation of the tolerance of the arm length (#5076)

  (done) Alignment of the Y green beam (#5084)

  (done) Preparation of beam dumps (#5047)

  (done) Health check of shadow sensors and the OSEM damping gain adjustment (#5061)

  (done) Alignment of the incident beam axis (#5073)

  (done) Loss measurement of the arm cavities (#5077)

  5088   Tue Aug 2 02:20:09 2011 kiwamuUpdateGreen LockingY arm green beam axis : done

I succeeded in locking the green light to the Y arm cavity, but it wasn't so robust. Something is unhealthy in the electronics.

I am leaving the Y green system as it is because I already can see a plenty of the green light flashing in the BS chamber.

So just a flashing of the green light is good enough for the in-vac-work.

DONE.

Quote from #5084

The next step is to refine some electronics in the PDH loops to get the green light locked to the Y arm cavity. 

 

  5087   Mon Aug 1 23:29:24 2011 Manuel, IshwitaUpdateWienerFilteringGetting Data by matlab

We tried to acquire data from the seismometers and the mode cleaner using the Matlab function

datalist = NDS2_GetData({'C1:PEM-SEIS_GUR1_X_IN1_DQ'}, 996258376 , 10, CONFIG.nds.C)

and encountered the following error

Warning: daq_request_data failed
 
??? Error using ==> NDS2_GetData
Fatal Error getting channel data.

The same error was obtained with the following other channels

C1:PEM-SEIS_GUR2_X_IN1_DQ

C1:PEM-SEIS_STS_1_X_IN1_DQ

But we are able to get data from channel

C1:LSC-MC_OUT_DQ

for the same gps time.

We checked with Dataviewer that the data are saved (we viewed data of last 24h) for every channel.

  5086   Mon Aug 1 23:26:32 2011 KojiUpdateLSCREFL33 PD

Old MZ PD (InGaAs 2mm, @29.5MHz) has been modified for REFL33.
There has been no choice for the 11MHz notch other than putting it on the RF preamp
as the notch in parellel to the diode eats the RF transimpedance at 33MHz.

I wait for judgement of Rana if the notch at the MAX4107 feedback is acceptable or not.

P8011390.JPG

Attachment 1: REFL33_schematic_110801_KA.pdf
REFL33_schematic_110801_KA.pdf
  5085   Mon Aug 1 22:30:59 2011 kiwamuSummaryGeneralWeekly report

Summary of the week ending July 31st.  Number of elog entries = 53


- SUS
 + ETMY-LR sensor looked strange. Something wrong.
 + Responses from the DC alignment bias to the shadow sensors and the oplevs were checked.
   --> ETMY shows the response with the opposite sign. Wired.
 + ETMY shadow sensors were examined in terms of the spectra.
   --> WF, AA and ADC noise looks reasonably low and not high enough to explain the low frequency noise.
 + Adjusted all the OSEM gains

- LSC
 + MICH noise budget is ongoing. WF filter needs to be greater than 21 dB to have dark noise of the PD greater than ADC noise
 + The arms became lockable

- CDS
 + modified and re-ran activateDQ.py.
 + c1iscex crashed for unknown reasons and we physically rebooted it.

- ALS (Fiber experiments)
 + PMC trans is sampled for the fiber beat-note measurement on the PSL table.
 + The beat-note signal between PSL and X end laser were obtained.
 + Some optics in the ETMY table were rearragend to have the Y green light aligned.

- ACS / ASS
 + incident beam axis has changed a lot.
 + X arm and Y arm ASS were reactivated.
  ---> The sign of some of the control gains had been wrong.
 + The incident beam axis and X/Y arm were re-aligned

- IOO/MCWFS
 + Some medm screens fixed.
 + Adjustment of the demodulation phase on each quadrant on WFS1 and WFS2 are done.
 + The sensing matrix (from optics to WFS sensors) were measured.

- OAF
 + c1pem was modified
 + plugged a seismometers to ADC through an AA board.
   --> channels are coming to the digital land
 
- Preparation for the invac work
 + 7 pieces of beam traps are available
 + Tolerance of the arm length is estimated to be +/- 2 cm.


- PSL/RefCav
 + ABSL is injected into the reference cavity. some flashing happened but no locking.
 + eddited the psl.db file to set EGUF and EGUL
 + turned RefCav heater and servo back on

  5084   Mon Aug 1 20:21:05 2011 kiwamuUpdateGreen LockingY arm green beam axis : aligned

The beam axis of the Y green light has been aligned.

Now I can see TEM00 mode is flashing on the ETMY camera.

 

-- (What I will do tonight)

The next step is to refine some electronics in the PDH loops to get the green light locked to the Y arm cavity.

If the beam isn't locked, I guess the in-vac-work will be so difficult because of the low intensity of the green light.

According to a brief check on the circuits, a low pass filter after the demodulation mixer is in a sad situation.

It doesn't pass any signals and in fact it behaves more like an absorber.

On the other hand, the modulation system looks fine to me because I was able to see the 270 kHz sideband converted into AM due to the fringing.

Quote from #5078

  (not yet) Alignment of the Y green beam (#5066)

  5083   Mon Aug 1 17:47:37 2011 steveUpdateASCwhat is not working

SUS-ETMY_QPD is not responding. It is reading zero in dataviewer and 4,400 counts on QPD MEDM screen.......must be wrong cable connected

IP-POS is sick. Last time alive 7-19-2011

IP-ANG beam is clipping on pick-up mirror at ETMY chamber. This will have to be fixed at the vent. The qpd itself is responding to light.

  5082   Mon Aug 1 16:06:05 2011 steveUpdateSUSoplevs are centered

Kiwamu, Koji and Steve,

Arm aligned separatly and oplev qpds were centered  including BS. Than we realigned manually DRM and centered their qpds.

ALL SUS oplevs were centered to resonating cavities.

  5081   Mon Aug 1 11:46:56 2011 ranaSummaryLSCTolerance of Arm length = 2 cm

wow. This Monte-Carlo matrix is one of the most advanced optical modeling things I have ever seen. We never had this for any of the interferometers before.

  5080   Mon Aug 1 08:52:37 2011 steveUpdateVACpreparation to vent

Both arms locked easely around 1V transmited.  We should recenter oplevs.

Attachment 1: pztsteering.png
pztsteering.png
Attachment 2: vac1Aug2011.png
vac1Aug2011.png
Attachment 3: PEM200d.jpg
PEM200d.jpg
  5079   Mon Aug 1 04:08:24 2011 kiwamuUpdateABSLArm length measurement : cavity kick technique

I made some attempts to measure the current length of the arm cavities by using the mass-kicking technique.

However unfortunately I am running out my energy to complete the measurement,

so I will finish the measurement at some time today.

I still have to set an appropriate kick amplitude. Right now I am injecting AWG into ETMY_LSC_EXC at 0.2 Hz with amplutde of 400 cnts.

I guess it needs a little bit more amplitude to get more psuedo-constant velocity.

Volunteers are always welcome !

 

(some notes)

The procedure was well-described in entry #555 by Dr.Stochino.

Here is just an example of the time series that I took today showing how the time series looks like.

ETMY_kick.png

  5078   Sun Jul 31 22:48:35 2011 kiwamuSummaryGeneralpreparation of the vent : status update

Status update for the vent preparation:

The punchline is : We can not open the chamber on Monday !

 

##### Task List for the vent preparation #####

  (not yet) Low power MC

  (not yet) Measurement of the arm lengths

  (not yet) Alignment of the Y green beam (#5066)

  (not yet) Measurement of the MC spot positions

  (80% done) Estimation of the tolerance of the arm length (#5076)

  (done) Preparation of beam dumps (#5047)

  (done) Health check of shadow sensors and the OSEM damping gain adjustment (#5061)

  (done) Alignment of the incident beam axis (#5073)

  (done) Loss measurement of the arm cavities (#5077)

Quote from #5048

Quote:

The vent will start from 1 st of August ! 

 

  5077   Sun Jul 31 22:35:35 2011 kiwamuUpdateLSCarm loss measurement : done

I did the measurement of the arm loss on both X and Y arm by running the armLoss script.

The results will be posted later.

Quote from #5074

I will measure the loss on the X and Y arm cavity tomorrow.

 

  5076   Sun Jul 31 17:28:34 2011 kiwamuSummaryLSCTolerance of Arm length = 2 cm

 Required arm length = 37.7974 +/- 0.02 [m]

This is a preliminary result of the estimation of the Arm length tolerance.

This number was obtained from a simulation based on Optickle.
Note that the simulation was done by considering misplacements in only the arm lengths while keeping PRCL, SRCL and MICH at the ideal lengths.
Therefore the tolerance will be somewhat tighter if misplacements in the central part are taken into account.

Next : check 3f signals, and include misplacements in PRCL, SRCL and MICH.



(Background)
We will re-position the ETMY/Y suspensions to adjust the arm lenghts during the coming vent.
To get a reasonable sensing matrix for LSC, the arm length must be adjisted within a certain precision.
So we need to know the tolerance of the arm lengths.
 

(How to estimate)
Optickle, a frequency domiain interferomtere simulator, is used to model the response of the 40m interferometer.
I buit a 40m model in Optickle, and in this model every optical distance is adjusted to the perfect length.
Then some offsets are added on the macroscopic position of ETMs to see what will happen in the LSC sensing matrix.
When putting the offsets, the amount of offsets are randomly assigned with a Gaussian distribution (see Figure.1).
Therefore the calculation is a Monte-Calro style, but this doesn't have to be a Monte-Calro
because the parameter space is only 2-dimensions (i.e. X-arm and Y-arm length) and it can be done by simply scanning the 2-dimentional parameter space.
The reason why the Monte-Calro style was chosen is because I wanted to expand this simulation
to a more general simulation which can handle PRCL, SRCL and MICH misplacements as well.
This time I ran the Monte-Calro 1000 times.
 
random2cm.png
Figure.1 History of random walk in X-Y arm lengths parameter space.
The position of ETMY and ETMX are randomly chosen with a Gaussian distribution function in every simulation.
This example was generated when \sigma_x = \sigma_y = 2 cm, where \sigma is the standard deviation of 
the Gaussian function. The number of simulation is 1000 times.
 
 
 

(Criteria)
I made two criteria for the acceptable sensing matrix as follows :
  (1) The decrease in the optical gain of the important signals (diagonal signals) must be within a factor of 3 (factor of ~ 0.5 in log scale).
  (2) MICH and SRCL signals are separated within a range of 60 - 120 deg in their demodulation phases on POP55.
 

(Results1 : sensing matrix)
Figure.2 shows the resultant sensing matrix with histograms when \sigma_x = \sigma_y = 2,
where \sigma_x, \sigma_y are the given standard deviation in the position of ETMX and ETMY.
The diagonal signals (in red-rectangular window) shows variation in their optical gain within a factor of 0.5 in log scale (factor of 3 in linear scale).
This satisfies my requirement (1) mentioned in the last section.
 
 
 
armsensMAT.png
Figure.2  A sensing matrix of the 40-m DRFPMI while changing the position of ETMX/Y by \sigma = 2 cm.
For convenience,  only REFL11, AS55, POP11 and POP55 are shown. They are the designed signal ports that
mentioned in the aLIGO LSC document (T1000298). In all the histograms, x-axis represents the optical gain in log scale in units of [W/m].
The y-axis is the number of events. The diagonal ports are surrounded by red rectangular window.
 
 
 
 
(Results2 : demodulation phase of MICH and SRCL on POP55)
Now a special attention should be payed on the MICH and SRCL signals on POP55.
Since MICH and SRCL are designed to be taken from POP55, they must be nicely separated in their demodulation phases.
Therefore the demodulation phase of MICH and SRCL has to be carefully examined.
The plot in Figure.3 is the resultant phase difference between MICH and SRCL on POP55 when \sigma_x = \sigma_y = 2 cm.
As shown in the plot the phase are always within a range of 60 - 120 deg, which satisfies my requirement (2) mentioned in the last section.
 
 
 
 
POP55phase2cm.png
 Figure.3 Difference in the demodulation phase of MICH and SRCL on POP55.
x-axis is the difference in the demodulation phase of MICH and SRCL, and y-axis the number of events.
 
 

(Notes on the Optickle model)
Optickle that I used is the one downloaded from the MIT CVS server and I believe this is the latest version.
In my current simulation I omitted some foldng mirrors including PR3, SR2 and SR3.
If those mirrors are added on the model, loss from those mirrors will affect the build up powers in all the cavities and hence changes the sensinag matrix somewhat.
I assumed that each optic has loss of 50 ppm in its HR surface.
Input power, after the MC, of 1 W is assumed.
The modulation depth are all 0.1 rad for 11MHz and 55MHz.
The model files were uploaded on the MIT CVS server and files reside under /export/cvs/iscmodeling/40m/fullIFO_Optickle.
More information about the CVS server can be found on aLIGO wiki.
 
  5075   Sun Jul 31 00:37:57 2011 kiwamuUpdatePSLRe : PSL table work

I think you made a simple mistake in your diagram -- the mixer must be replaced by a summer circuit. Otherwise you cannot do the PDH lock.

Quote from #5069

LIGO_block_diagram.png

  5074   Sun Jul 31 00:05:52 2011 kiwamuUpdateLSCscript for loss measurement : modified

I modified the script armloss so that the channel names in the script are properly adopted to the new CDS.

Additionally I disabled the ETMX(Y)_tickle command in the script.

The tickle command puts some offsets on the LSC signal to let the arms pass through a fringe until it gets locked, but apparently we don't need it because the arms are loud enough.

A brief check showed that the script ran fine.

I will measure the loss on the X and Y arm cavity tomorrow.

Quote from #5067

Next : Health-check for the X arm ASS, the loss measurements.

 

  5073   Sat Jul 30 21:04:23 2011 kiwamuUpdateASCY arm ASS fixed

The X arm ASS was also fixed. So both X and Y arm ASS are now back to normal.

Now we can align the arms any time from the buttons on the C1IFO_CONFIGURE window.

 

(notes)

The reason why the servo didn't work was that the sign of some control gains had been flipped.

This was exactly the same situation as that in the Y arm ASS (#5067).

Quote from #5067

The servos of C1ASS for the Y arm and the beam axis alignments were fixed.

  5072   Sat Jul 30 20:41:50 2011 ranaUpdatePSLRefCav Stabilization back on

 Untitled.png

I turned the RefCav heater and servo back on a couple days ago. At first it was stabilizing again at a low setpoint, but in reality the right temperature (~40 C). After fixing the in-loop signal offsets, the setpoint now correctly reflects the actual temperature.

Jenny is going to calibrate the sensors using some kind of dunking cannister next week.

  5071   Sat Jul 30 19:06:25 2011 ryan, ranaUpdatePSLReturn of the PSL temperature box

Quote:

The PSL temperature box has returned to service, with some circuit modifications. The 1k resistors on all the temp. sensor inputs (R3, R4, R7, R8, R12, R12) were changed to 0 Ohm. Also, the 10k resistors R26, R28, R29, and R30 were changed to 10.2k metal film. The DCC document will be updated shortly. There is now an offset in the MINCOMEAS channel compared to the others, which will be corrected in the morning after looking at the overnight trend.

 Forgot to do this in May. Have just changed the values in the psl.db file now as well as updating them live via Probe.

To make the appropriate change, I took the measured offset (5.31 deg) and added 2x this to the EGUF and EGUL field for the MINCO_MEAS channel. (see instructions here)

Committed the .db file to the SVN.

attached plot shows 8 days of trend with 5.31 degC added to the black trace using the XMGRACE Data Set Transformations

Attachment 1: rctempbox.png
rctempbox.png
  5070   Sat Jul 30 10:03:32 2011 JennyUpdateComputer Scripts / ProgramsMode matching

I ended up having to switch to a different mode-matching solution, because I was unable to find the f = 572.7 mm lens. See my next elog entry (5069).

  5069   Sat Jul 30 10:01:35 2011 JennyUpdatePSLPSL table work

I've been working on the PSL table to put together a setup so that I can measure the reference cavity's response to a temperature step increase at the can surrounding it. My first step was to mode match the beam coming from the AP table to the cavity.

I implemented my mode matching solution. I ended up using a different one from the one I last elogged about. Here is the solution I used:

Two lenses: f = 1016.7.6 mm at -0.96 m and f = 687.5 mm at -0.658 m. (I set my origin at the polarizing beam splitter--the spot where I want my beam to match the beam coming from the PMC, so all waists are behind that point). Below is what it should look like.

modematchpic.pngmodematchinfo.png

What I did on the table:

  • Before placing lenses I aligned the beam and added a 1/2-wave plate between the two polarizing beam splitters to change the polarization of the beam from S to P.
  • I aligned the beam so that it reflected off of the cavity opening (monitoring the reflected power with a photodetector connected to an oscilloscope and tweaking the alignment to maximize the reflected signal). 
  • I then placed the lenses at -0.93 and -0.64 mm because the exact spots were blocked by optics being used in another setup.
  • I reasoned that since the fitting for the initial waist is so uncertain, the lens position being off by a few cm will not produce the dominating source of error. I am now driving the laser frequency using a lock-in as a function generator to drive the laser temperature at ~1 Hz. I'm then monitoring the power transmitted by the reference cavity with a camera connected to a TV monitor. I will use this setup to improve my mode matching.

Here's a picture of the PSL table with the lenses and mirror I added. The beam is redirected by a mirror and then a polarizing beam splitter. Past the beam splitter is another lens (f=286.5 mm), which was already in place from the mode matching of the beam from the PMC to the reference cavity.

modematch_setup_pic.png

Here is a block diagram of my intended experimental setup:

LIGO_block_diagram.png

I am going to try to lock the laser to the cavity given my preliminary mode matching and then go back and improve it later. My next step is to find a frequency range for dithering the voltage sent to the PZT. To do this I will:

  • Measure the transfer function (amplitude response) of the PZT using a photodiode. The power outputted by the laser varies with driving frequency.
  • Find a frequency region in which the amplitude response is low.
Attachment 3: LIGO_block_diagram.png
LIGO_block_diagram.png
  5068   Sat Jul 30 07:07:28 2011 kiwamuUpdateASCMCASS status

Since we will measure (and hopefully adjust) the spot positions on the MC suspensions prior to the vent, MCASS is necessary for it.

 

#######################

Here is the MCASS status so far:

  + Valera worked on MCASS on the last February, and basically no progress after he left.

  + The MCASS model had been completed in C1IOO.mdl.

  + He made some useful scripts, including mcassup, mcassOn/Off, senseMCdecenter, senseMCmirrro and senseMCdofs.

       Summary of those scripts can be found in his entry #4355.

  + We haven't closed the MCASS loops.

  + The control filters are still blank.

  + We haven't put any elements on the input and output matrices.

  + Some parameters for the dithering oscillators and demodulation systems were properly set.

     So we can get the demodulated signals by simply running mcassUp and mcassOn. (This essentially corresponds to the A2L measurement.)

  + The PIT motions are driven at 10, 11 and 12 Hz for MC1, 2 and 3 respectively. For YAW, the frequencies were chosen to be 11.5, 12.5 and 13.5 Hz.

  + Some medm windows were prepared but not as refined as that of ASS.

  + Valera performed a measurement of the spot positions by using MCASS. The results are summarized in #4660.

  + We made an estimation about the beam clearance on the Faraday based on the measured spot positions (#4674)

 

#########################

So, it seems we should be able to at least measure the spot positions soon by using his scripts.

  5067   Sat Jul 30 06:24:06 2011 kiwamuUpdateASCY arm ASS fixed

The servos of C1ASS for the Y arm and the beam axis alignments were fixed.

Now we can correctly run the Y arm ASS from the C1IFO_CONFIGURE window as usual.

The sign of some control gains had been flipped for some reasons, so I changed them to the correct signs.

 

Next : Health-check for the X arm ASS, the loss measurements.

Quote from #5062

I tried using C1ASS to align the incident beam and suspensions on the Y arm, but it didn't work.

 

  5066   Sat Jul 30 05:11:45 2011 SureshUpdateGreen LockingY- end table clean-up

The optics on the Y-end table which required to be moved have been repositioned.  Please see the attached pic for details.

The green beam is not yet aligned to the cavity. That is my next task.

Y-end_table_work.png

  5065   Sat Jul 30 02:47:43 2011 ranaUpdatePSLABSL Laser crystal temp left largely excited & left unattended for more than 3hours

 Hmm. Should have only been +/- 1 GHz. Some setting got changed apparently...

This is a part of the RefCav temperature measurement setup. You'll get an elog from Jenny very soon.

  5064   Sat Jul 30 00:33:33 2011 KojiUpdatePSLABSL Laser crystal temp left largely excited & left unattended for more than 3hours

I found that the ref cav trans CCD view was blinking with 30-50 fringe amplitudes. This meant the laser freq was swinging ~50GHz.

I checked the ABSL laser and the SG out of a lock-in amplifier was connected to the slow input.

This was shaking the laser temp from 29degC to 46degC. This was the cause of the fringe swinging.
This big excitation changing the output power too as the temp was changed across it mode-hop region.

I have disconnected the excitation from the laser no matter how useful experiments were took place as there was no e-log entry about this.

I need the explanations

1. Why our precious laser is exposed to such a large swing of temperature?

2. Why the excitation is left like that without any attendance?

3. Why there was no elogging about this activity?

  5063   Fri Jul 29 18:43:02 2011 Manuel, IshwitaUpdatePEMplugging seismometers to ADC

[Manuel, Ishwita, Jenne, Jamie]

We changed the C1PEM model and the names of the C1:PEM channels.

We reinstalled the blue breakout box, since the purple one still didn't work.

So, now the AA board channels are connected as follows...

C1 = C1:PEM-SEIS_GUR1_X

C2 = C1:PEM-SEIS_GUR1_Y

C3 = C1:PEM-SEIS_GUR1_Z

C4 = C1:PEM-SEIS_GUR2_X

C5 = C1:PEM-SEIS_GUR2_Y

C6 = C1:PEM-SEIS_GUR2_Z

C7 = C1:PEM-SEIS_STS_1_X

C8 = C1:PEM-SEIS_STS_1_Y

C9 = C1:PEM-SEIS_STS_1_Z

C11 = C1:PEM-SEIS_STS_2_X

C12 = C1:PEM-SEIS_STS_2_Y

C13 = C1:PEM-SEIS_STS_2_Z

C14 = C1:PEM-SEIS_STS_3_X

C15 = C1:PEM-SEIS_STS_3_Y

C16 = C1:PEM-SEIS_STS_3_Z

C17 = C1:PEM-ACC_MC1_X

C18 = C1:PEM-ACC_MC1_Y

C19 = C1:PEM-ACC_MC1_Z

C20 = C1:PEM-ACC_MC2_X

C21 = C1:PEM-ACC_MC2_Y

C22 = C1:PEM-ACC_MC2_Z

Although the channels for all 3 STS-2 seismometers are made but only one is installed. So only Channels C1 to C9 are now in operation...

We checked the data from the plugged channels with the Dataviewer. We could see the peak whenever someone jumped in the lab. Even Kiwamu jumped and saw his signal.

  5062   Fri Jul 29 16:25:06 2011 kiwamuUpdateASCbeam axis and Y arm aligned

Last night I aligned the incident beam axis and the Yarm by touching the PZT mirrors and the suspensions.

I didn't estimate how good they were aligned, but I guess the Y arm is now ready for the Y green light.

 Next : Y green alignment and the MC spots measurement / alignment.

 

 ++ Motivation ++

Prior to the coming vent we want to have the Y arm, incident beam axis and Y green light aligned so that we can align some necessary optics in the chamber.

Also alignment of the incident beam will allow us to re-position the incident beam alignment monitor (i.e. IPPOS and IPANG).

Our plan was to first align the Y arm using the ASS system and then align the Y green light to the Y arm.

 

++ what I failed ++

First I was trying to measure the spot positions on the MC mirrors to make me sure the beam axis has/hasn't changed.

Also I was going to align the MC suspensions to have nice spot position on each suspension using the MCASS system

because this will help us checking the beam clearance in the Faraday and perhaps re-positioning of the Faraday during the coming vent.

But essentially I failed and eventually gave up because MCASS didn't work. It seems that MCASS needs some modifications in the scripts.

Then, to make me feel better I moved on to the Y arm and beam axis alignment.

 

++ what I did ++

I tried using C1ASS to align the incident beam and suspensions on the Y arm, but it didn't work.

However the drive signals from ASS and its demodulated signals looked fine. Only the feedback did not work correctly.

Every time I enabled the feedback paths, the arm just lost the lock. Something is wrong in the feedback paths.

Then I started to align the cavity by my hands while looking at the demodulated signal from each LOCKIN module.

I aligned the things until each demodulated signal fluctuates around zero.

At the end the beam spots on the ETMY and ITMY camera looked well-aligned and the transmitted light became larger by a factor of 2ish.

  5061   Fri Jul 29 16:04:51 2011 steveUpdateSUSETMX, PRM & SRM OSEM gains set

Jamei fixed the computer. Now I had a healthier ETMX with SIDE signal that allowed me to set gains to get Q of 5

The Y arm restore locked the arm at TRY -out 2.4V,    the X arm still did not lock or align to lock. It has some fringes.

ETMX  POS gain   22  -> 17,              PIT  3  ->  1,                    YAW  3,                     SIDE     -170    ->    -110

PRM                        50  ->  30,                     1,                                       2,                                       50

SRM                        55  ->  20,                    2.4  ->  1.5,                      4.8  ->  2,                          140   ->    100

Now all SUS OSEM gains are set

  5060   Fri Jul 29 12:39:26 2011 jamieUpdateCDSc1iscex mysteriously crashed

c1iscex was behaving very strangely this morning.  Steve earlier reported that he was having trouble pulling up some channels from the c1scx model.  I went to investigate and noticed that indeed some channels were not responding.

While I was in the middle of poking around, c1iscex stopped responding altogether, and became completely unresponsive.  I walked down there and did a hard reset.  Once it rebooted, and I did a burt restore from early this morning, everything appeared to be working again.

The fact that problems were showing up before the machine crashed worries me.  I'll try to investigate more this afternoon.

  5059   Fri Jul 29 12:25:54 2011 Ishwita, ManuelUpdatePEMSTS-2 seismometer box

The 'Bacardi' STS-2 seismometer was tested with the "purple" breakout box and it was found out that all the three axes gave a voltage of 11 V (as shown on the screen of the oscilloscope) before pressing the auto-zero button and after pressing it the voltage shown was 6 V. We tried again the blue box and it was working perfectly after pushing the auto zero button (the auto zero took a few seconds). The power of the purple box is still on, we will wait a few hours, to see if something changes.

  5058   Thu Jul 28 21:52:40 2011 ranaUpdateComputersanother attempt to use pianosa
  1. Pianosa doesn't cache the SVN pwds so you need to re-enter at each SVN up or commit. This is different from the rest of our workstations. We need to determine what behavior we want.
  2. Tried to use the netGPIB scripts:

pianosa:gpib 0> ./readSR785.csh rb2
rb2
netgpibdata.py: Command not found.

  5057   Thu Jul 28 19:49:12 2011 SonaliUpdatePSLPMC trans beam aligned.

Kiwamu and I aligned the PMC transmitted beam the incident beam going to PMC today.

I learnt how to lock the PMC using the digital controls.

  5056   Thu Jul 28 17:04:04 2011 NicoleUpdateSUSPhotosensor Head Calibration Curve

Here is the calibration curve (displacement versus voltage output) for the photosensor head that I made with the S5971 photodiodes and L9337 LEDs. This was made using a regular mirror. The linear region appears to be between 0.4 and 0.75cm. I will need to arrange the photosensor head so it measures displacements in the linear region of this plot. This plot was made using a 287 ohm resistor.

187K_resistor.jpg

  5055   Thu Jul 28 16:26:42 2011 steveUpdateSUSITMX, ITMY & ETMY OSEM gains adjusted

 

OSEM damping gains were adjusted by observing  real time dataviewer to get Q of 5

OSEMs were kicked up one by one with 200 counts  ~1sec. The error signal was optimized to get 1/2 of exitation amplitude at the 5th sinusoid wave.

C1: SUS-ITMX_SUSPOS_N1 gain   111      ->        65,             PIT        7.2      ->    8,           YAW        12      ->     6,          SIDE     280

                  ITMY                                      277       ->     120,                        19.2     ->     7,                             24     ->     19,                        420    ->      470

                  ETMY                                       10        ->       32,                          20      ->     3,                             20     ->     10,                          50

                  ETMX                                       22        ->      25,                             3,                                                3,                                              -170  

ETMX having problems:  1, YAW can not be excited

                                             2, SIDE has no error signal in dataviewer. Sensing voltage on MEDM screen 0.142V

 

                      

 

  5054   Thu Jul 28 16:10:34 2011 kiwamuUpdateLSCBoth arm locked

[Nicole / Jamie / Rana / Kiwamu]

  The X arm and Y arm have been locked.

The settings for the locking were stored on the usual IFO_CONFIGURE scripts, so anybody can lock the arms.

In addition to that Nicole, Jamie and Rana re-centered the beam spot on the ETMY_TRANS camera and the TRY PD.

The next step is to activate the C1ASS servo and align the both arms and beam axis.

 

 

Xarm locking notes:

* Changed TRX gain from -1 to -0.02. Without this 50x reduction the arm power was not normalized.

* Had to fix trigger matrix to use TRX for XARM and TRY for YARM. Before it was crazy and senseless.

* Lots of PZT alignment. It was off by lots.

* Yarm trans beam was clipping on the steering mirrors. Re-aligned. Needs to be iterated again. Be careful when bumping around the ETMY table.

* YARM gain was set to -2 instead of -0.2. Because the gain was too high the alignment didn't work right.

ALWAYS HAVE an OPEN DATAVIEWER with the standard ARM channels going when doing ANY INTERFEROMETER WORK.

THIS IS THE LAW.

  5053   Thu Jul 28 16:00:28 2011 kiwamuUpdateComputer Scripts / Programsanother offset script : offset2

A new offset-zeroing script has been developed and it is ready to run.

 

 The motivation is to replace the old zeroing script called offset by a better one because this old script somehow failed to revert the gain settings on a given filter bank.

The new script, named offset2, does the same job, but uses tdsavg instead of using ezcaservo. So it doesn't screw up the gain settings.

Additionally the structure of the script is much simpler than the old offset script, and fewer ezca-functions.

I will modify some scripts which use the old offset script so that all the offset-zeroing is done by offset2.

 

P.S.

Useful scripts are listed on the 40m wiki

http://blue.ligo-wa.caltech.edu:8000/40m/Computers_and_Scripts/All_Scripts

  5052   Thu Jul 28 13:51:00 2011 SonaliUpdateGreen LockingZHL-32A-S.

Initially I was using RFPD-1611to get the IR beat frequency. Its gain was not very high, so I was getting a very low signal of power -37 dBm.

I used ZHL-32A-S with a gain of 25 dBm to amplify it before feeding it into the spectrum analyser.

I connected the ground of the amplifier circuit to the red of the power supply, which blew the amplifier.

I learnt that there is a small tab indicating the ground side of the BNC to banana connectors which I should have noticed.

I learnt to plug in the side with th little tab on it into the ground of the power supply. (Learnt it the hard way I guess!!)

 

 

  5051   Thu Jul 28 02:33:04 2011 JenneUpdateLockingYarm flashing, but not yet locked

Because I'm too lazy to write a cohenrent elog right now, here's my notes that I wrote while working tonight:

Elog notes, 27July2011

Aligned Xarm, just to check on it.  Had to flip sign of TRX in DCPD filter bank (to gain of -1) to make the signal positive.

Restored Yarm, see some slight flashing, but no lock yet.
Adjusted phase rotation of AS55 from 56.5deg to 60deg, just by-eye trying to maximize AS55I, my arm error signal. AS55I goes from ~ -40 to +60 counts

Tried fitzing with Yarm gain, flipping sign, incr gain. No real change in signals, or flashing.


Incr. ETMY oplev gains to -0.4 from -0.2
Engaged ELP35's on Pit and Yaw, to be more similar to other optics.  However, right now all of the optics have different things in their filter banks.  Why??

Arm is flashing pretty reliably now, but still not locking.  The trigger threshold is always satisfied, so that's not it.

ELOG V3.1.3-